# Chapter 0 Hadron Production Processes

### v1

@@ 1 Interacting Hadrons
The quark-gluon configurations displayed in Fig. 8 are building blocks of the QCD core and could be taken as a basis for extended studies. The core eigenstates are obtained by diagonalization within the quark-gluon configuration space, which is demanding but not prohibitive. The result is a spectrum of energy-sharp, discrete eigenstates because of confinement. CQM and bag models of hadron structure are coming closest to that kind of picture. However, for describing physical hadrons realistically, it has to be taken into account that these model states are embedded in the continuum of hadronic scattering states [^12] In molecular, atomic, and nuclear physics such states are known as bound states in the continuum (BIC)..
As theoretical constructs, the bare quark-gluon states correspond to closed channels with inseparable constituents under normal environmental conditions. The (multi-)hadron-hadron scattering states are open channels with asymptotically free particles. Depending on the case, they are composed either of pure meson and baryon-antibaryon for $B=0$ systems, or of mixed meson-baryon scattering states in $B\neq 0$ objects. If there would be 6-quark objects, e.g., the notorious and experimentally still searched for $S=-2$ H-dibaryon hypothesized by Jaffe [77], baryon-baryon channels would be involved as well. The observation of the non-strange $d^{\ast}(2380)$ dibaryon, probably a double-$\Delta^{0}$ configuration, was reported by the WASA-COSY collaboration [78]. Theoretical studies find rather compact d-baryon wave functions [79]. In either case, the set of discrete channel quantum numbers, defined by the coupled irreducible representations of the hadronic constituents, must coincide with those of the core [^13] It’s worth noting that the best known and stable dibaryon is the deuteron, although not of the compact stature as expected for exotic states.. Functional methods and LQCD will in principle account for such effects, as the good reproduction of the experimental mass spectrum in Fig. 9 shows.
The basic mechanisms governing the interplay of quark-gluon and hadron-hadron components of hadron wave functions are easily understood in Hamiltonian formulation (named after William Rowan Hamilton (1805-1865)). The description of physical hadrons demands solving a coupled channels problem.
@@ 2 Interactions of QCD Core and Hadron Scattering Configurations
Consider a QCD core state $|Q\rangle$ of bare mass $M_{Q}$ which is coupled by the interactions $V_{QC}=V^{\dagger}_{CQ}$ to a set of open hadron scattering channels $C$ of total channel mass $M_{C}=\sum_{h\in C}m_{h}$, given by the sum over the masses $m_{h}$ of all asymptotically free hadrons $h=p,n,\pi\ldots$. Standard projection methods allow us to recast the set of coupled equations by a fully equivalent dispersive, non-Hermitian polarization self-energy operator [^14] An operator $A$ is Hermitian if $A=A^{\dagger}=(A^{T})^{*}=(A^{*})^{T}$ does not change under transposition ($T$) and complex conjugation (*). A real number $x$ is Hermitian, a complex number $z=x+iy$ in non-Hermitian. Any non-Hermitian operator can be decomposed in a Hermitian and an anti-Hermitian operator (named after Charles Hermite (1822-1901)). acting on the quark-gluon core:
@@
$\Delta_{Q}(w_{\alpha})=\sum_{C}\int\frac{d^{3}q}{(2\pi)^{3}}c|V_{QC}(\mathbf{q})|^{2}\frac{\mathcal{P}}{w_{\alpha}-w_{C}(q)}\>$ (4)
@@
[…]
Including the self-energy, i.e., the coupling to the open hadron channels, the bear (model) states $|Q\rangle$ are transformed to the polarized - in this model the physical - resonances $|R\rangle$. They are determined by a generalized, dispersion eigenvalue problem. The spectrum of resonances is defined by the (complex) roots $\mathcal{M}_{R}$ of the equation
@@
A closer inspection of the dispersion equations reveals that if the QCD configuration $|Q\rangle$ interacts with $N^{(Q)}_{C}$ channel states $|C\rangle$, Eq. (9) has in total $N_{R}=N^{(Q)}_{C}+1$ solutions of discrete complex eigenvalues $\mathcal{M}_{R_{n}}$. Hence, for each fundamental QCD configuration, the channel coupling will produce a spectrum of $N^{(Q)}_{C}$ satellite states $|R_{n}\rangle$, $n\leq N^{(Q)}_{C}$ plus a solution corresponding to the unperturbed state $|Q\rangle$ of mass eigenvalue $\mathcal{M}_{R_{n}}=M_{R_{n}}-\frac{i}{2}\Gamma_{R_{n}}$. The spectroscopic strength of $|Q\rangle$ is distributed over the spectrum of core-like eigenstates. Typically, but in detail depending on the channel spectrum and the interactions $V_{CQ}$, the state of energy closest to $M_{Q}$ will acquire the largest spectroscopic factor.
In order to maintain analyticity, theoretical approaches should account for the energy dependence of all parts of the self-energies. However, for obvious reasons, little to nothing is known about the intrinsic structures of self-energies because their full knowledge is equivalent to the exact solution of the complete CC problem. In practice, typically the width, i.e., the imaginary part of the channel self-energy, is approximated by functional forms of the proper behavior at the threshold. Close to the threshold of a partial wave with orbital angular momentum $\ell$ and channel momentum $\mathbf{q}$, the coupling matrix elements behave as $V_{CQ}(|\mathbf{q}|\sim q^{\ell}$, implying for the width $\Gamma_{CQ}\sim q^{2\ell}$. Widely used parameterizations are the so-called Blatt-Weisskopf form factors (named after John Markus Blatt (1921-1990) and Victor Frederick “Wiki” Weisskopf (1908-2002)) $F_{\ell}(q/q_{0})$ [80]. These form factors account for the centrifugal barrier effects which determine the behavior of an outgoing scattering wave at the interaction radius. They are given by spherical Hankel functions (named after Hermann Hankel (1839- 1873). In [81], they were generalized to hadron resonance scattering. Explicit formulas and practical applications are discussed, e.g., in Ref. [38].
Once the imaginary part of a self-energy is fixed, the real part is obtained consistently by dispersion theory. The Chew-Mandelstam method (named after Geoffrey Foucar Chew (1924-2019) and Stanley Mandelstam (1928-2016)) accomplishes that goal by evaluating a once-subtraction dispersion integral where the widths are parameterized by Blatt-Weisskopf form factors. In the SAID approach, these effects are incorporated [82].
@@ 3 Polarization Self-energies in Hadron Scattering Amplitudes
where $\Gamma_{Q}=\sum_{C}|\rho_{C}\langle V_{CQ}\rangle|^{2}$ and $x_{CQ}=\rho_{C}|\langle V_{CQ}\rangle|^{2}/\Gamma_{Q}$ are the branching ratios, to be evaluated at the appropriate channel momenta. As a consequence, the scattering amplitude of channel $C$ becomes the sum of the scattering amplitudes from the generic channel interactions and the amplitude produced by $\Sigma_{C}$. The latter gives rise to resonances which interfere with the weakly energy-dependent in-channel amplitudes. As a function of energy, the respective partial wave cross sections will show peak structures on top of a smooth background. Occasionally, the resonances are accompanied by an interference pattern which may lead to irregular line shapes.
The scattering channels will interact also among themselves. In combination with kinematical and structural conditions, e.g., openings of particle emission thresholds or energy-dependent accidental cancellations or enhancements among resonance and background components, the channel-channel interaction amplitudes may produce resonance-like structures in the partial wave cross section which, however, being of purely dynamical origin, will fall out of the systematics expected by symmetries and the related selection rules. Care has to be taken to identify, isolate, and separate model-dependent effects which are achieved by comparing independently derived results of different approaches. Careful studies of the evolution of the amplitudes in the complex plane by Argand diagrams (named after Jean-Robert Argand (1768-1822)) [83] will reveal the nature of the structure. Dalitz plots, introduced in 1953 by Richard Henry Dalitz (1925-2006) while studying kaon decay [84], are indispensable tools for investigations of three-body reaction channels by revealing correlations between decay products.
@@ 4 Channel Coupling in Meson-Meson Scattering
The above formalism is a general scheme, applicable to any interacting quantum system at any scale. The same kind of basic mechanisms are acting in baryon-baryon, meson-baryon, and meson-meson systems. Prominent cases are the formation of pion-nucleon resonances like the first excited state of the nucleon given by the well-known Delta-resonance $P_{33}(1232)$ of width $\Gamma_{\Delta}=120~{}\mathrm{MeV}$, corresponding to a lifetime of about $t_{1/2}\sim 10^{-23}$ sec. The widths and finite lifetimes of the members of the $J^{P}=1^{-}$ vector meson octet point in the same direction. The isovector ($I=1$) $\rho(770)$ meson, for example, is embedded into the $\pi\pi$ $P$-wave continuum, resulting in a decay width of $\Gamma_{\rho}\sim 150~{}\mathrm{MeV}$ with a two-pion branching ratio of 99.9%. Isospin selection rules are affecting the rho-meson by excluding the $\rho^{0}\to\pi^{0}\pi^{0}$ decay channel.
@@
Occasionally, the $\rho$ meson is considered as the gauge boson of a broken hidden local chiral symmetry which, however, is distinct from the broken global chiral symmetry of which the pion is the respective Goldstone (named after Jeffrey Goldstone (born 1933)) boson (named after Satyendra Nath Bose) [85]. Moreover, the non-strange vector mesons $\rho,\omega,\phi$ play a central role in Sakurai’s Vector Meson Dominance (VMD) model (named after Jun John Sakurai (1933-1982)) [86] where they account for the (virtually admixed) hadronic part of the photon by which the photon couples to hadronic matter.
@@
The $\rho$- and $\omega$-mesons were detected first in 1961 at LBL. Remarkably, the $\omega(782)$ meson, being the isoscalar partner of the rho-meson, is a rather long-lived object with a small width of about $\Gamma_{\omega}\sim 8~{}\mathrm{MeV}$. The increase in life-time/decrease in decay width by about a factor of 20 is an effect of spin-parity conservation laws, inhibiting decays of an isoscalar vector meson into two pions. An even more extreme case is the $\phi(1020)$ meson which is the singlet partner of the $\omega$-meson. Although located in a mass region with many open channels, a width as small as $\Gamma_{\phi}=4.249\pm 0.013~{}\mathrm{MeV}$ [38] is observed. The long life-time of the $\phi$-meson is caused by the dominant $s\bar{s}$ structure which requires the creation of $u\bar{u}$ or $d\bar{d}$ pairs out of the vacuum before a decay into lighter mesons, e.g., the prevailing $\phi\to K\bar{K}$ process, can occur, as expressed in the OZI rule which is also responsible for extraordinarily small widths of $c\bar{c}$ states like $J/\psi$.
@@
Since presently dedicated meson production facilities are not available, $N+\pi\to N^{\prime}+2\pi$ and $N+\gamma\to N^{\prime}+2\pi$ reactions are used to investigate the otherwise inaccessible $\pi\Delta(1232)$ and $\rho(770)N$ final states, also aiming at separating these final states kinematically (see, for instance, Ref. [87]). Such indirect methods are also applied for extracting hadron-hadron cross sections from hadron production data in high-energy nucleon-nucleus and nucleus-nucleus collisions at the operating hadron facilities.
@@ 1 Overview
Modeling hadron production and especially the search for resonances on the quantitative level is a demanding task. Several groups have developed coupled channel theories, derived sophisticated models apt for practical work, and cast the theoretical results into numerical codes, ready for large-scale analyses of the data measured at the experimental sites and, not least, serving also for the preparation of experimental campaigns.
Figure 10: Diagram illustrating the Bethe-Salpeter equation of the T-matrix for scattering of a hadron or photon (dashed line) on a nucleons or any other baryon (full line). The bare interactions, i.e. the Born terms, are denoted by $V$, the scattering amplitude $T$ includes the full scattering series, summed to all orders. The evolution of the intermediate system is described by the Bethe-Salpeter propagator $\mathcal{G}_{BS}$ (see text).
The existing approaches are utilizing a covariant description of the scattering process. The major step is to construct out of an effective Lagrangian the interaction matrix elements [^15] Joseph-Louis Lagrange (1736-1813) worked out the formalism of analytical classical mechanics, and the variational methods which in quantized form became the essential tools of QFT. which serve as input to a system of coupled Bethe-Salpeter equations (named after Hans Albrecht Eduard Bethe (1906-2005) and Edwin Salpeter (1924-2008)) [^16] Hans Bethe received the Nobel Prize in Physics in 1967 for his work on the theory of stellar nucleosynthesis, which is illustrated in Fig. 10. Couplings between channels, e.g., the aforementioned case of $|KN\rangle\leftrightarrow|\pi\Sigma\rangle$ coupling and interference in $\Lambda(1405)$ spectroscopy, have to be taken into account. In Fig. 11, the elemental ingredients of coupled channels approaches for hadron production on the nucleon are displayed in diagrammatical form.
@@
where $V$ is the matrix containing the full set of elemental channel interactions. $w=\sqrt{s}$ is the available center-of-mass energy. $p$ and $p^{\prime}$ are the incoming (outgoing) hadron (on-shell) four-momenta. Intermediate propagation is described by the BS propagator $\mathcal{G}_{BS}$. The integration over the energy variable $q^{0}$ is performed in closed form by contour integration, leading to the reduced propagator $G_{BS}(\mathbf{q};\sqrt{s})$. That step in fact requires special attention because it involves also a projection to positive energy states, hence effectively eliminating the vacuum contributions. The price to pay is that interactions have to account implicitly for the eliminated degrees of freedom by effective, phenomenological coupling constants. A reduction scheme widely used, e.g., in NN scattering is the Blankenbecler-Sugar (BbS) approach (Richard Blankenbecler (born 1933) and Robert Sugar [88]). The BbS projection reduces the full BSE to a three-dimensional problem by preserving covariance and two- and three-body unitarity.
The solution of a large system of coupled integral equations defined in complex algebra is a formidable numerical task even for modern computing facilities. In order to optimize the numerical workload, the full CC problem is solved piecewise by constructing first the K-matrix and then retrieving the full T-matrix in a second step.
For that goal, the reduced propagator $G_{BS}$ is decomposed into its real and imaginary parts. Accordingly, the (reduced) BSE separates can be rewritten in terms of two (nested) equations. The scattering series generated by $Re(G_{BS})$ can be summed separately, unaffected by $Im(G_{BS})$ leading to the K-matrix equation
@@
which accounts for all off-shell contributions. The momentum integrals must be regularized, which is accomplished by attaching to the bare interaction $V$ vertex form factors:
@@
where $q^{2}$ denotes the four-momentum squared of the involved particle of mass $m$ which mediates the interactions.
@@
$T(\mathbf{p},\mathbf{p}^{\prime};w)=K(\mathbf{p},\mathbf{p}^{\prime};w)+i\int\frac{d^{3}q}{(2\pi)^{3}}T(\mathbf{p},\mathbf{q};w)\mathrm{Im}\left[G_{BS}(q;w)\right]K(\mathbf{a},\mathbf{p}^{\prime};w)\>.$ (14)
This is still a system of coupled integral equations but of a much simpler structure than before by the fact that the imaginary part of $G_{BS}$ is given by Dirac $\delta$-distributions in energies and/or powers of momenta, projecting the intermediate channels to their respective on-shell kinematics. The reduced BSE is finally solved by expanding the matrix elements into partial wave components, by which the remaining integration over the angles of the momenta in the chosen reference frame is performed analytically.
Denoting the partial wave components by their total spin $J$, orbital angular momentum $L$ and parity $P$, a further reduced system of coupled integral equations is obtained:
[…]
Final, initial, and intermediate $meson-baryon$ channels are denoted by $f,i,j$, respectively. Spectral distributions of the intermediate configurations are taken into account by integration over the respective mass distribution $A_{j}(\mu_{j}$, thus allowing a detailed description of unstable states like $f_{0}\sigma(500),\rho(770)\ldots$ attached to a stable baryon or the propagation and interaction of decaying baryons like $\Delta_{33}(1232)$
For stable particles of mass $m_{j}$, the mass distributions reduce to $A_{j}=\delta(\mu_{j}-m_{j})$. In the other cases, the integrals are evaluated by numerical integration formulas, thus replacing the integrals by a sum over a discrete set of mesh points. Then, the integral equations reduce to a system of coupled linear equations that are solved numerically. The essence of such a treatment is that unstable states are represented by a discrete distribution of states with fractional spectroscopic strengths as defined by the value of the spectral distribution at the mesh points and the integration weights, see, e.g.[89]. Unitarity is fulfilled as long as $V$ is Hermitian.

### v2

@@ 1 Interacting Hadrons
The quark-gluon configurations displayed in Fig. 8 are the building blocks of the QCD core and could be taken as a basis for extended studies. The core eigenstates are obtained by diagonalization within the quark-gluon configuration space, which is demanding but not prohibitive. The result is a spectrum of energy-sharp, discrete eigenstates due to confinement. CQM and bag models of hadron structure are coming closest to that kind of picture. However, for describing physical hadrons realistically, it must be taken into account that these model states are embedded in the continuum of hadronic scattering states [^13] In molecular, atomic, and nuclear physics, such states are known as bound states in the continuum (BIC)..
As theoretical constructs, the bare quark-gluon states correspond to closed channels with inseparable constituents under normal environmental conditions. The (multi-)hadron-hadron scattering states are open channels with asymptotically free particles. Depending on the case, they are composed either of pure mesons and baryon-antibaryon pairs for $B=0$ systems or of mixed meson-baryon scattering states in $B\neq 0$ objects. If there would be 6-quark objects, e.g., the notorious and experimentally still searched for S=2S=-2 H-dibaryon hypothesized by Jaffe [88], baryon-baryon channels would be involved as well. The observation of the non-strange $d^{\ast}(2380)$ dibaryon, probably a double-$\Delta^{0}$ configuration, was reported by the WASA-COSY collaboration [89]. Theoretical studies find rather compact d-baryon wave functions [90]. In either case, the set of discrete channel quantum numbers, defined by the coupled irreducible representations of the hadronic constituents, must coincide with those of the core [^14] It’s worth noting that the best known and stable dibaryon is the deuteron, although it is not of the compact stature expected for exotic states.. Functional methods and LQCD will, in principle, account for such effects, as the good reproduction of the experimental mass spectrum in Fig. 9 shows.
The basic mechanisms governing the interplay of quark-gluon and hadron-hadron components of hadron wave functions are easily understood in the Hamiltonian formulation (named after William Rowan Hamilton (1805-1865)). The description of physical hadrons demands solving a coupled-channel problem.
@@ 2 Interactions of QCD Core and Hadron Scattering Configurations
Consider a QCD core state $|Q\rangle$ of bare mass $M_{Q}$ which is coupled by the interactions $V_{QC}=V^{\dagger}_{CQ}$ to a set of open hadron scattering channels $C$ of total channel mass $M_{C}=\sum_{h\in C}m_{h}$, given by the sum over the masses $m_{h}$ of all asymptotically free hadrons $h=p,n,\pi\ldots$. Standard projection methods allow us to recast the set of coupled equations by a fully equivalent dispersive, non-Hermitian polarization self-energy operator [^15] An operator $A$ is Hermitian if $A=A^{\dagger}=(A^{T})^{\ast}=(A^{\ast})^{T}$ does not change under transposition ($T$) and complex conjugation ($\ast$). A real number $x$ is Hermitian, a complex number $z=x+iy$ is non-Hermitian. Any non-Hermitian operator can be decomposed into a Hermitian and an anti-Hermitian operator (named after Charles Hermite (1822-1901)). acting on the quark-gluon core:
@@
$\Delta_{Q}(w_{\alpha})=\sum_{C}\int\mathcal{P}\frac{d^{3}q}{(2\pi)^{3}}\frac{|V_{QC}(\mathbf{q})|^{2}}{w_{\alpha}-w_{C}(q)}\>$ (4)
@@
$\langle|V_{CQ}(q_{\alpha C})|^{2}\rangle=\frac{1}{4\pi}\int d\hat{\mathbf{q}}|V_{CQ}(\mathbf{q})|^{2}_{|q=q_{\alpha C}}\>.$ (8)
Including the self-energy, i.e., the coupling to the open hadron channels, the bare (model) states $|Q\rangle$ are transformed into the polarized - in this model the physical - resonances $|R\rangle$. They are determined by a generalized, dispersion eigenvalue problem. The spectrum of resonances is defined by the (complex) roots $\mathcal{M}_{R}$ of the equation.
@@
A closer inspection of the dispersion equations reveals that if the QCD configuration $|Q\rangle$ interacts with the $N^{(Q)}_{C}$ channel states $|C\rangle$, Eq. (9) has a total of $N_{R}=N^{(Q)}_{C}+1$ solutions of discrete complex eigenvalues $\mathcal{M}_{R_{n}}$. Hence, for each fundamental QCD configuration, the channel coupling will produce a spectrum of $N^{(Q)}_{C}$ satellite states $|R_{n}\rangle$, $n\leq N^{(Q)}_{C}$ plus a solution corresponding to the unperturbed state $|Q\rangle$ of mass eigenvalue $\mathcal{M}_{R_{n}}=M_{R_{n}}-\frac{i}{2}\Gamma_{R_{n}}$. The spectroscopic strength of $|Q\rangle$ is distributed over the spectrum of core-like eigenstates. Typically, but in detail depending on the channel spectrum and the interactions $V_{CQ}$, the state of energy closest to $M_{Q}$ will acquire the largest spectroscopic factor.
To maintain analyticity, theoretical approaches should account for the energy dependence of all parts of the self-energies. However, for obvious reasons, little to nothing is known about the intrinsic structures of self-energies because their full knowledge is equivalent to the exact solution of the complete CC problem. In practice, typically the width, i.e., the imaginary part of the channel self-energy is approximated by functional forms of the proper behavior at the threshold. Close to the threshold of a partial wave with orbital angular momentum $\ell$ and channel momentum $\mathbf{q}$, the coupling matrix elements behave as $V_{CQ}(|\mathbf{q}|)\sim q^{\ell}$, implying a width of $\Gamma_{CQ}\sim q^{2\ell+1}$. Widely used parameterizations are the so-called Blatt-Weisskopf form factors (named after John Markus Blatt (1921-1990) and Victor Frederick Weisskopf (1908-2002) $F_{\ell}(q/q_{0})$ [91]. These form factors account for the centrifugal barrier effects, which determine the behavior of an outgoing scattering wave at the interaction radius. They are given by spherical Hankel functions (named after Hermann Hankel (1839- 1873). In [92], they were generalized to hadron resonance scattering. Explicit formulas and practical applications are discussed, e.g., in Ref. [41].
Once the imaginary part of a self-energy is fixed, the real part is consistently obtained through dispersion theory. The Chew-Mandelstam method (named after Geoffrey Foucar Chew (1924-2019) and Stanley Mandelstam (1928-2016)) accomplishes that goal by evaluating a once-subtraction dispersion integral, where the widths are parameterized by Blatt-Weisskopf form factors. In the SAID approach, these effects are incorporated [93].
@@ 3 Polarization Self-energies in Hadron Scattering Amplitudes
where $\Gamma_{Q}=\sum_{C}|\rho_{C}\langle V_{CQ}\rangle|^{2}$ and $x_{CQ}=\rho_{C}|\langle V_{CQ}\rangle|^{2}/\Gamma_{Q}$ are the branching ratios, to be evaluated at the appropriate channel momenta. As a consequence, the scattering amplitude of channel $C$ becomes the sum of the scattering amplitudes from the generic channel interactions and the amplitude produced by $\Sigma_{C}$. The latter gives rise to resonances that interfere with the weakly energy-dependent in-channel amplitudes. As a function of energy, the respective partial wave cross sections will show peak structures on top of a smooth background. Occasionally, the resonances are accompanied by an interference pattern, which may lead to irregular line shapes.
The scattering channels will also interact among themselves. In combination with kinematical and structural conditions, e.g., openings of particle emission thresholds or energy-dependent accidental cancellations or enhancements among resonance and background components, the channel-channel interaction amplitudes may produce resonance-like structures in the partial wave cross section which, however, being of purely dynamical origin, will fall outside the systematics expected by symmetries and the related selection rules. Care must be taken to identify, isolate, and separate model-dependent effects, which are achieved by comparing independently derived results from different approaches. Careful studies of the evolution of the amplitudes in the complex plane by Argand diagrams (named after Jean-Robert Argand (1768-1822)) [94] will reveal the nature of the structure. Dalitz plots, introduced in 1953 by Richard Henry Dalitz (1925-2006) while studying kaon decay [95], are indispensable tools for investigations of three-body reaction channels by revealing correlations between decay products.
@@ 4 Channel Coupling in Meson-Meson Scattering
The above formalism is a general scheme, applicable to any interacting quantum system at any scale. The same types of basic mechanisms are acting in baryon-baryon, meson-baryon, and meson-meson systems. Prominent cases include the formation of pion-nucleon resonances like the first excited state of the nucleon represented by the well-known Delta-resonance $P_{33}(1232)$ of width $\Gamma_{\Delta}=120~\mathrm{MeV}$, corresponding to a lifetime of about $t_{1/2}\sim 10^{-23}$ sec. The widths and finite lifetimes of the members of the $J^{P}=1^{-}$ vector meson octet indicate the same trend. The isovector ($I=1$) $\rho(770)$ meson, for example, is embedded in the $\pi\pi$ $P$-wave continuum, resulting in a decay width of $\Gamma_{\rho}\sim 150~\mathrm{MeV}$ with a two-pion branching ratio of 99.9%. Isospin selection rules affect the rho-meson by excluding the $\rho^{0}\to\pi^{0}\pi^{0}$ decay channel.
@@
Occasionally, the $\rho$ meson is considered to be the gauge boson of a broken hidden local chiral symmetry which, however, is distinct from the broken global chiral symmetry of which the pion is the respective Goldstone (named after Jeffrey Goldstone (born 1933)) boson (named after Satyendra Nath Bose) [96]. Moreover, the non-strange vector mesons $\rho,\omega,\phi$ play a central role in Sakurai’s Vector Meson Dominance (VMD) model (named after Jun John Sakurai (1933-1982)) [97] where they account for the (virtually admixed) hadronic part of the photon, through which the photon couples to hadronic matter.
@@
The $\rho$- and $\omega$-mesons were first detected in 1961 at LBL [98]. Remarkably, the $\omega(782)$ meson, being the isoscalar partner of the $\rho$-meson, is a rather long-lived object with a small width of about $\Gamma_{\omega}\sim 8~\mathrm{MeV}$. The increase in lifetime/decrease in decay width by about a factor of 20 is an effect of isospin conservation. Inhibiting the decay of an isoscalar vector meson into two pions. An even more extreme case is the $\phi(1020)$ meson, which is the singlet partner of the $\omega$-meson. Although located in a mass region with many open channels, a width as small as $\Gamma_{\phi}=4.249\pm 0.013~\mathrm{MeV}$ [41] is observed. The long life-time of the $\phi$-meson is caused by the dominant $s\bar{s}$ structure which requires the creation of $u\bar{u}$ or $d\bar{d}$ pairs out of the vacuum before a decay into lighter mesons, e.g., the prevailing $\phi\to K\bar{K}$ process, can occur, as expressed in the OZI rule which is also responsible for extraordinarily small widths of $c\bar{c}$ states like $J/\psi$.
@@
Since currently dedicated meson production facilities are not available, $N\pi\to N^{\prime}2\pi$ and $N\gamma\to N^{\prime}2\pi$ reactions are used to investigate the otherwise inaccessible $\pi\Delta(1232)$ and $\rho(770)N$ final states, also aiming to separate these final states kinematically (see, for instance, Ref. [99]). Such indirect methods are also applied to extract hadron-hadron cross sections from hadron production data in high-energy nucleon-nucleus and nucleus-nucleus collisions at the operating hadron facilities.
@@ 1 Overview
Modeling hadron production, and especially the search for resonances on a quantitative level, is a demanding task. Several groups have developed coupled channel theories, derived sophisticated models suitable for practical work, and converted the theoretical results into numerical codes, ready for large-scale analysis of the data measured at the experimental sites and, not least, also serving to prepare for experimental campaigns. A detailed review of these ground-breaking developments is given elsewhere in this encyclopedia [100].
Figure 10: Diagram illustrating the Bethe-Salpeter equation of the T-matrix for scattering of a hadron or photon (dashed line) on a nucleon or any other baryon (full line). The bare interactions, i.e. the Born terms, are denoted by $V$; the scattering amplitude $T$ includes the full scattering series, summed to all orders. The evolution of the intermediate system is described by the Bethe-Salpeter propagator $\mathcal{G}_{BS}$ (see text).
The existing approaches utilize a covariant description of the scattering process. The major step is to construct, from an effective Lagrangian, the interaction matrix elements [^16] Joseph-Louis Lagrange (1736-1813) developed the formalism of analytical classical mechanics and the variational methods, which, in quantized form, became the essential tools of QFT. These serve as input to a system of coupled Bethe-Salpeter equations (named after Hans Albrecht Eduard Bethe (1906-2005) and Edwin Salpeter (1924-2008)) [^17] Hans Bethe received the Nobel Prize in Physics in 1967 for his work on the theory of stellar nucleosynthesis, which is illustrated in Fig. 10. Couplings between channels, e.g., the aforementioned case of $|KN\rangle\leftrightarrow|\pi\Sigma\rangle$ coupling and interference in $\Lambda(1405)$ spectroscopy, must be taken into account. In Fig. 11, the elemental ingredients of coupled channel approaches for hadron production on the nucleon are displayed in diagrammatic form.
@@
where $V$ is the matrix containing the full set of elemental channel interactions. $w=\sqrt{s}$ is the available center-of-mass energy. $p$ and $p^{\prime}$ are the incoming (outgoing) hadron (on-shell) four-momenta. Intermediate propagation is described by the BS propagator $\mathcal{G}_{BS}$. The integration over the energy variable $q^{0}$ is performed in closed form by contour integration, leading to the reduced propagator $G_{BS}(\mathbf{q};\sqrt{s})$. That step, in fact, requires special attention because it involves a projection to positive energy states, hence effectively eliminating the vacuum contributions. The price to pay is that interactions must implicitly account for the eliminated degrees of freedom by effective, phenomenological coupling constants. A reduction scheme widely used, e.g., in NN scattering, is the Blankenbecler-Sugar (BbS) approach (Richard Blankenbecler (born 1933) and Robert Sugar [101]). The BbS projection reduces the full BSE to a three-dimensional problem by preserving covariance and two- and three-body unitarity.
The solution of a large system of coupled integral equations defined in complex algebra is a formidable numerical task, even for modern computing facilities. In order to optimize the numerical workload, the full CC problem is solved piecewise by first constructing the K-matrix and then retrieving the full T-matrix in a second step.
[…]
@@
which accounts for all off-shell contributions. The momentum integrals must be regularized, which is accomplished by attaching to the bare interaction $V$ vertex form factors. A widely used, but not unique choice is
@@
where $q^{2}$ denotes the four-momentum squared of the involved particle of mass $m$, which mediates the interactions.
@@
$T(\mathbf{p},\mathbf{p}^{\prime};w)=K(\mathbf{p},\mathbf{p}^{\prime};w)+i\int\frac{d^{3}q}{(2\pi)^{3}}T(\mathbf{p},\mathbf{q};w)\mathrm{Im}\left[G_{BS}(q;w)\right]K(\mathbf{q},\mathbf{p}^{\prime};w)\>.$ (14)
This is still a system of coupled integral equations, but it has a much simpler structure than before because the imaginary part of $G_{BS}$ is given by Dirac $\delta$-distributions in energies and/or powers of momenta, projecting the intermediate channels to their respective on-shell kinematics. The reduced BSE is finally solved by expanding the matrix elements into partial wave components, through which the remaining integration over the angles of the momenta in the chosen reference frame is performed analytically.
Denoting the partial wave components by their total spin $J$, orbital angular momentum $L$, and parity $P$, a further reduced system of coupled integral equations is obtained:
$T^{JLP}_{f\,i}(w)=K^{JLP}_{f\,i}(w)+{\rm i}\sum_{j}\int^{\infty}_{\mu_{j0}}d\mu_{j}K^{JLP}_{j\,i}(\mu_{j},w)A_{j}(\mu_{j};w)T^{JLP}_{f\,j}(\mu_{j};w)\>.$ (15)
Final, initial, and intermediate mesonbaryonmeson-baryon channels are denoted by $f,i,j$, respectively. Spectral distributions of the intermediate configurations are taken into account by integrating over the respective mass distribution $A_{j}(\mu_{j})$, thus allowing a detailed description of unstable states like $f_{0}$ or $\sigma(500)$, $\rho(770)\ldots$ attached to a stable baryon, or the propagation and interaction of decaying baryons such as $\Delta_{33}(1232)$.
For stable particles of mass $m_{j}$, the mass distributions reduce to $A_{j}=\delta(\mu_{j}-m_{j})$. In the other cases, the integrals are evaluated using numerical integration formulas, thus replacing the integrals with a sum over a discrete set of mesh points. Then, the integral equations reduce to a system of coupled linear equations that are solved numerically. The essence of such a treatment is that unstable states are represented by a discrete distribution of states with fractional spectroscopic strengths, as defined by the value of the spectral distribution at the mesh points and the integration weights, see e.g.[102]. Unitarity is fulfilled as long as $V$ is Hermitian.